RC meeting and Run Plan, Tue Oct. 9

From HallCWiki
Jump to: navigation, search
  • Status
    • Data taken so far [1]
    • Beam quality [2]

  • Issues: Resolved
    • HMS dipole LN2 JT4 valve
    • SHMS rotation: checked by Steve, until we have the breaks working this [3] may happen; be very careful when rotating.

  • Requiring attention
    • The Hall C truck ramp beacon [4]
    • Monitor HALLC:p [5]
    • Change to SHMS angle of 6.65 deg: if spokespeople would like to go lower than that without an access (software limit is 6.48 deg per rotation gui), try to arrange with Steve L. for assistance with rotation to this angle (it would have to happen during the day)

  • Plans - Physics
    • Take data at Q2=3. W=3.14 x=0.25 theta_Shms=12.42 to reach 65% of desired stat. goal.
    • Change SHMS angle to 9.42 deg and take LH2 and Al data